# Fully Spiking Neural Network for Legged Robots

### v1

@@ Abstract
In recent years, legged robots based on deep reinforcement learning have made remarkable progress. Quadruped robots have demonstrated the ability to complete challenging tasks in complex environments and have been deployed in real-world scenarios to assist humans. Simultaneously, bipedal and humanoid robots have achieved breakthroughs in various demanding tasks. Current reinforcement learning methods can utilize diverse robot bodies and historical information to perform actions. However, prior research has not emphasized the speed and energy consumption of network inference, as well as the biological significance of the neural networks themselves. Most of the networks employed are traditional artificial neural networks that utilize multilayer perceptrons (MLP). In this paper, we successfully apply a novel Spiking Neural Network (SNN) to process legged robots, achieving outstanding results across a range of simulated terrains. SNN holds a natural advantage over traditional neural networks in terms of inference speed and energy consumption, and their pulse-form processing of body perception signals offers improved biological interpretability. To the best of our knowledge, this is the first work to implement SNN in legged robots.
@@ I Introduction
The increasing adoption of mobile robots, which are equipped with continuous high-dimensional observations and action space, to tackle a wide range of intricate tasks in real-world situations emphasizes the critical importance of exceptional control algorithms. Currently, the limited onboard energy resources of most robots pose a significant challenge, as this constraint hinders their ability to operate continuously and cost-effectively. Consequently, there is an immediate demand for developing energy-efficient solutions for the seamless control of these autonomous machines. Deep reinforcement learning (DRL) employs deep neural networks (DNNs) as potent function approximators for learning optimal control strategies for intricate tasks [1, 2], through directly mapping the original state space to the action space [3, 4]. Nonetheless, the remarkable performance of DRL frequently comes at the expense of substantial energy consumption and slower execution speeds, making them unsuitable for various applications. Additionally, the execution speed of control strategies employing DNNs tends to be slower in comparison to the operational speed of the motion units. This discrepancy often results in a step-like behavior in the control signals, causing negative impacts on the performance of the system.
@@
  • For the first time, we have implemented SNNs on a policy network in the legged robot simulated in Isaac Gym[21], enabling the encoding of each dimension of the observation and action space within a single population of neurons using a learnable acceptance domain. Furthermore, we have successfully integrated this method and achieved successful training outcomes using techniques such as imitation learning and trajectory history.
  • We have decreased the decimation time to allow for more frequent updates during the learning process. Numerous experiments have consistently shown that SNNs perform exceptionally well even at ultra-high control frequencies.
  • We achieved successful training on three typical legged robots: quadruped robotic dog a1, bipedal robot Cassie, and MIT humanoid, which fully demonstrate the feasibility and superiority of our method for application on legged robots (Fig. 1).
@@ II-A Reinforcement Learning for Legged Robotics Locomotion
In recent years, there has been a rapid development of reinforcement learning techniques in the realm of achieving fast, stable, and adaptive locomotion for legged robots. [22] bridges the gap between reinforcement learning simulation and actual reality by modeling the actuator through neural networks. [23] utilizes the teacher-student training framework to integrate environmental parameters with preconceptions which enables quadruped robots to adapt to complex terrains. [24] introduces a large number of auxiliary rewards and gait parameter control to implement multiple gaits for a single policy. This work brings a lot of inspiration to the design of quadruped rewards. [25] combines reinforcement learning with computer vision to complete simultaneous locomotion and manipulation. Unlike traditional reinforcement learning, Generative adversarial imitation learning (GAIL) is an approach for imitating behaviors from reference datasets by integrating a generative adversarial network. Based on the GAIL, Adversarial Motion Priors (AMP) combines the task reward and imitation reward to make the agent complete the tasks based on the action being similar to the reference dataset. For learning unlabeled references dataset, [26] uses a skill discriminator additionally and enables quadruped to distinguish and master multiple gaits, and perform backflips not found in the dataset. [27] combines RMA and AMP, allowing the quadruped with the ability to traverse challenging terrains rapidly. However, the above methods are realized by ANNs. Therefore, it cannot take into account the high-frequency and energy-saving advantages of SNNs.
Figure 2: Firstly, the observations are encoded by the encoder as $n$ independent distributions, uniformly distributed over the range of the observation. These encoded distributions are then further processed by the population, resulting in the generation of the corresponding spikes. Then the neurons in the input populations are responsible for encoding each dimension of the observation and they drive a multi-layered and fully connected SNN. During the forward timesteps, the activities of each output population in PopSAN are decoded to determine the corresponding action dimension. This means that the neural network takes in observations, processes them through the SNN, and then decodes the resulting activities to generate the appropriate action for the given situation.
@@ III SNNs based Locomotion in Isaacgym
We chose to train and test the performance of our algorithm on Isaac Gym[21]. Isaac Gym is a simulation platform designed for robotics. It offers realistic physics simulation, specialized support for legged robots, integration with NVIDIA technologies, customization options, and an active community.
@@ III-A SNN based Policy Network
We utilize a population-coded spiking actor-network (PopSAN)[37] that is trained in conjunction with a deep critic network using the DRL algorithms (including RMA and AMP). During training, the PopSAN generated an action $\alpha$in $\mathbb{R}^{N}$ for a given observation, $s$, and the deep critic network predicted the associated state value $V$($s$) or action-value $Q$($s$, $\alpha$), which in turn optimized the PopSAN, in accordance with a chosen DRL method (Fig. 2). In the PopSAN architecture, the encoder module encodes each dimension of the observation into the activity of a specific neuron population. During forward propagation, these input populations stimulate a multi-layer fully-connected SNN. The SNN then produces activity patterns in the output populations. At the end of every $T$ timesteps, these activity patterns are decoded to determine the corresponding action dimensions (as outlined in Algorithm 1).
@@
To construct the SNN, the current-based leaky-integrate-and-fire (LIF) model of a spiking neuron is employed. This model is utilized for building the SNN architecture. The dynamics of the LIF neurons are controlled by a two-step model, as elaborated in Algorithm 1: i) integrating the presynaptic spikes $o$ into current $c$; and ii) integrating the current $c$ into membrane voltage $v$; $d_{c}$ and $d_{v}$ are the current and voltage decay factors. In this implementation, a neuron fires a spike when its membrane potential surpasses a certain threshold. We adopted the hard-reset model, which means that upon spiking, the membrane potential is instantly reset to the resting potential. The resulting spikes are then transmitted to the post-synaptic neurons within the same inference timestep, assuming zero propagation delay. This approach enables efficient and synchronous transmission of information within the SNN.
@@
Figure 3: RMA is a system composed of two subsystems: the base policy $\pi$ and the adaptation module $\phi$. The training of RMA consists of two phases. Training the Base Policy (Phase 1): In the first phase, the base policy $\pi$ is trained using PopSAN. It takes as input the current state $x_{t}$, the previous action $\alpha_{t-1}$, and the privileged environmental factors $e_{t}$. These environmental factors are encoded into a latent extrinsics vector $z_{t}$ using the environmental factor encoder $\mu$. Training the Adaptation Module (Phase 2):In the second phase, the adaptation module $\phi$ is trained to predict the extrinsics $\widehat{z_{t}}$ based on the history of states and actions. This training is done using supervised learning with on-policy data. The adaptation module learns to capture the relationship between the state-action history and the corresponding extrinsics. By training the base policy and the adaptation module in these two phases, RMA is able to learn and adapt to the environment in a more effective manner.
Next, we combined the superiority of snn with the recently advanced algorithm RMA. Figure 3 illustrates that RMA system comprises two interconnected subsystems: the base policy $\pi$ and the adaptation module $\phi$, which collaborate harmoniously to facilitate continuous adaptation in a wide variety of environmental setups, thereby enabling smooth and uninterrupted online operations. The base policy is trained through reinforcement learning in simulation, utilizing privileged information about the environment configuration $e_{t}$, which includes factors like friction, payload, and more. By leveraging the knowledge of the vector $e_{t}$, the base policy can effectively adapt to the specific characteristics of the given environment. The process begins by encoding the environment configuration vector $e_{t}$ into a latent feature space $z_{t}$ through an encoder network $\mu$. This latent vector, referred to as the extrinsics, is then combined with the current state $x_{t}$ and the previous action $\alpha_{t-1}$ as inputs to the base policy. The base policy then generates predictions for the desired joint positions of the robot, denoted as $a_{t}$. The policy $\pi$ and the environmental factor encoder $\mu$ undergo a collaborative training process using RL within a simulated environment.
@@
Regrettably, direct deployment of this policy is not feasible in the real world due to the unavailability of $e_{t}$. To overcome this challenge, we need to estimate the extrinsics during runtime, a task performed by the adaptation module $\phi$. The key insight here is that when we instruct the robot joints to perform a specific movement, the actual movement executed deviates from the intended movement, and this deviation is influenced by the extrinsics. Rather than relying on privileged information, we can leverage the recent history of the agent’s state to estimate the extrinsics vector. Specifically, the purpose of $\phi$ is to estimate the extrinsics vector $z_{t}$ solely based on the recent state and action history of the robot, without any direct access to $e_{t}$. At training time, since both the state history and the extrinsics vector $z_{t}$ can be computed in simulation, we can train this module using supervised learning techniques.
@@
In addition, we have successfully combined SNN with AMP and achieved similar performance to ANN on legged robots. Figure 4 provides a schematic overview of the system. The motion dataset $M$ consists of a collection of reference motions, where each motion $m^{i}$ $=$ $\widehat{q}_{t}^{i}$ is represented as a sequence of poses $\widehat{q}_{t}^{i}$. The motion clips can be obtained from a variety of sources, including motion capture (mocap) recordings of real-life actors or artist-authored keyframe animations. The motion of the simulated robot is controlled by a policy $\pi(\alpha_{t}|s_{t},g)$ that maps the state of the character $s_{t}$ and a given goal $g$ to a distribution over actions $\alpha_{t}$. The actions generated by the policy dictate the desired target positions for proportional-derivative (PD) controllers, which are located at each joint of the robot. These controllers then generate control forces that propel the robot’s motion in accordance with the specified target positions. The goal $g$ specifies a task reward function $r_{t}^{G}=r^{G}(s_{t},\alpha_{t},s_{t+1},g)$, which defines high-level objectives that the robot must fulfill. These objectives can include tasks such as walking in a specific direction or executing a punch towards a designated target. The style objective $r_{t}^{S}=r^{S}(s_{t},s_{t+1})$ is specified by an adversarial discriminator, trained to discern between motions captured in the dataset and motions generated by the robot itself. The style objective serves as a task-agnostic motion prior, offering an a-priori estimate of the naturalness or style of a given motion, regardless of the specific task. By doing so, the style objective motivates the policy to generate motions that closely resemble the behaviors observed in the dataset.
@@ III-B Training of Legged Robot using SNN on Isaacgym
The SNN parameters are updated using the extended spatiotemporal backpropagation introduced in [38]. We used the rectangular function $z(v)$, defined in [39], to approximate the gradient of a spike. The gradient of the loss with respect to the SNN parameters for each layer $k$ are computed by collecting the gradients backpropagated from all the timesteps:
@@ IV Experiments
The goals of our experiments are the followings: i) To validate the feasibility of SNNs on robots with high-latitude complex environments and complex dynamics models. ii) Verify the advantages of SNNs over ANNs in terms of ultra-high flatness control. iii) Identify rewards that SNNs excel on rl tasks and make improvements on poorly performing rewards. We evaluated our method on the Isaac Gym with accurate physics simulation and flexible and customizable framework for robot models and environments.
We mainly tested the performance of the following robots: A1, Cassie and MIT Humanoid where each snn is trained within $1,500,000$ iterations until convergence. The final visualization and metrics such as base velocity x, y, yaw, mean episode length and position torque are used for evaluation.
@@ IV-A Simulation Setup
To create a diverse range of environments, we imported several URDF files from recent works. These files include models such as A1, Anymal-b, Anymal-c, Cassie, MIT Humanoid, and more. Once imported, we leverage the capabilities of the Isaac Gym simulator to create new environments for each of these models. To add variety, we make use of the inbuilt fractal terrain generator provided by the simulator. This terrain generator enables us to generate different types of terrain, ranging from plain terrains to uneven terrains with various topographical features. Our policy operates at a control frequency of up to 500Hz, due to our SNN-based approach. This allows us to achieve high-frequency control, resulting in fast and accurate adjustments to the system. In Table I listed a comprehensive list of environmental variations and their ranges.
@@ IV-B Performances of High Frequency Control using SNNs
We conducted tests on the robots mentioned above specifically for linear and angular velocity tracking tasks. In these tests, we considered tracking velocity (linear and angular) as the primary positive rewards. Additionally, we incorporated penalties for base height that is too low, excessive acceleration, and instances where the robot falls, etc.
@@
For A1, we conducted training and testing in several terrain environments, including pyramid stairs like terraces (upstairs and downstairs), pyramids with sloping surfaces, hilly terrains, terrains with discrete obstacles, and terrains covered with stepping stones. On the other hand, Cassie is solely trained in a trapezoidal pyramid environment and MIT Humanoid in a plain terrain.
To further investigate the advantages of SNNs in high-frequency control scenarios, we deliberately increased the simulation environment’s time step (dt) to 2.5 times the original default ANNs task, resulting in a frequency of 500Hz. By comparing the performance of SNN with ANN under these conditions, our aim is to establish whether SNN outperforms ANN in real-world environment that require high-frequency control.
@@
The rationale behind this investigation stems from the fact that robots using ANN control algorithms often face limitations due to their onboard energy constraints. As a result, these robots can typically achieve only a policy inference frequency of 100Hz, which falls significantly behind the execution frequency of the motors. Conversely, SNN offers the potential to enhance the quality of policy inference and deliver more real-time control performance due to its energy-efficient nature and deployment on specialized processors. If SNN can achieve similar or even superior performance compared to ANN under high-frequency control, it would indicate the superiority of SNN in real-world environments.
In Figure 5, the effectiveness of our method under high-frequency control is clearly demonstrated. The performance of the A1 robot in tracking velocity x is highly commendable, showcasing its ability to accurately follow the desired trajectory. Moreover, when considering the impact of the intricate terrain environment, the fluctuation range of the linear velocity in the y-axis direction remains within an acceptable range. This further highlights the robustness of our method in tackling challenging terrains.
Furthermore, the results obtained in the yaw-axis angular velocity tracking are also notable. Although occasional spikes occur, indicating moments when the robot deviates from the existing policy framework, these instances can be attributed to the robot’s inclination to explore alternative approaches when faced with difficult terrain. This adaptive behavior showcases the robot’s ability to go beyond pre-existing policies and dynamically adapt its actions to overcome obstacles.

### v2

@@ Abstract
In recent years, legged robots based on deep reinforcement learning have made remarkable progress. Quadruped robots have demonstrated the ability to complete challenging tasks in complex environments and have been deployed in real-world scenarios to assist humans. Simultaneously, bipedal and humanoid robots have achieved breakthroughs in various demanding tasks. Current reinforcement learning methods can utilize diverse robot bodies and historical information to perform actions. However, prior research has not emphasized the speed and energy consumption of network inference, as well as the biological significance of the neural networks themselves. Most of the networks employed are traditional artificial neural networks that utilize multilayer perceptrons (MLP). In this paper, we successfully apply a novel Spiking Neural Network (SNN) to process legged robots, achieving outstanding results across a range of simulated terrains. SNN holds a natural advantage over traditional neural networks in terms of inference speed and energy consumption, and their pulse-form processing of body perception signals offers improved biological interpretability. Applying more biomimetic neural networks to legged robots can further reduce the heat dissipation and structural burden caused by the high power consumption of neural networks. To the best of our knowledge, this is the first work to implement SNN in legged robots.
@@ I Introduction
The increasing adoption of mobile robots, which are equipped with continuous high-dimensional observations and action space, to tackle a wide range of intricate tasks in real-world situations emphasizes the critical importance of exceptional control algorithms. Currently, the limited onboard energy resources of most robots pose a significant challenge, as this constraint hinders their ability to operate continuously and cost-effectively. Consequently, there is an immediate demand for developing energy-efficient solutions for the seamless control of these autonomous machines. Deep reinforcement learning (DRL) employs deep neural networks (DNNs) as potent function approximators for learning optimal control strategies for intricate tasks [1, 2], through directly mapping the original state space to the action space [3, 4]. Nonetheless, the remarkable performance of DRL frequently comes at the expense of substantial energy consumption and slower execution speeds, making them unsuitable for various applications. Additionally, the execution speed of control strategies employing DNNs tends to be slower in comparison to the operational speed of the motion units. This discrepancy often results in a step-like behavior in the control signals, causing negative impacts on the performance of the system. Moreover, the reliance of current Deep Neural Networks (DNNs) on high-power GPUs for computation significantly impacts the structural design of robots, necessitating the inclusion of substantial active cooling systems.
@@
  • For the first time, we have implemented SNNs on a policy network in various legged robots simulated in Isaac Gym[21], enabling the encoding of each dimension of the observation and action space within a single population of neurons using a learnable acceptance domain. Furthermore, we have successfully integrated this method and achieved successful training outcomes using techniques such as imitation learning and trajectory history.
  • Our approach presents a considerable advantage over conventional ANNs in terms of energy efficiency. This advantage holds substantial significance for enhancing the structural integrity and reducing the costs associated with robot development.
  • The research affirms the exceptional performance demonstrated by SNNs in the domain of high-frequency robot control, coupled with their significant edge over ANNs in attenuating signal noise, which enhances their robustness in practical situations.
@@ II-A Reinforcement Learning for Legged Robotics Locomotion
In recent years, there has been a rapid development of reinforcement learning techniques in the realm of achieving fast, stable, and adaptive locomotion for legged robots. [22] bridges the gap between reinforcement learning simulation and actual reality by modeling the actuator through neural networks. [23] utilizes the teacher-student training framework to integrate environmental parameters with preconceptions which enables quadruped robots to adapt to complex terrains. [24] introduces a large number of auxiliary rewards and gait parameter control to implement multiple gaits for a single policy. This work brings a lot of inspiration to the design of quadruped rewards. [25] combines reinforcement learning with computer vision to complete simultaneous locomotion and manipulation. Unlike traditional reinforcement learning, Generative adversarial imitation learning (GAIL) is an approach for imitating behaviors from reference datasets by integrating a generative adversarial network. Based on the GAIL, Adversarial Motion Priors (AMP) [26] combines the task reward and imitation reward to make the agent complete the tasks based on the action being similar to the reference dataset. For learning unlabeled references dataset, [27] uses a skill discriminator additionally and enables quadruped to distinguish and master multiple gaits, and perform backflips not found in the dataset. [28] combines RMA and AMP, allowing the quadruped with the ability to traverse challenging terrains rapidly. However, the above methods are realized by ANNs. Therefore, it cannot take into account the high-frequency and energy-saving advantages of SNNs.
Figure 2: The observations are initially encoded by the encoder as $n$ independent distributions that are uniformly distributed over the observation range. After encoding, the population processes the distributions, resulting in spike generation. The neurons in the input populations encode each observation dimension and drive a multi-layered, fully connected SNN. During forward timesteps in PopSAN, the activities of each output population are decoded to determine the corresponding action dimension. This implies that the neural network receives observations, processes them using the SNN, and decodes the resulting activities to determine the appropriate action for the specific situation.
@@ III SNNs based Locomotion in Isaacgym
We opted to evaluate the efficacy of our algorithm by training and testing it using Isaac Gym[21]. Isaac Gym is a simulation platform specifically created for robotics applications. The software provides a realistic physics simulation, specialized support for legged robots, integration with NVIDIA technologies, customizable features, and a vibrant community.
@@ III-A SNN based Policy Network
We employ a population-coded spiking actor-network (PopSAN)[38] that is trained in tandem with a deep critic network using the DRL algorithms (including RMA and AMP). During training, the PopSAN generated an action $\alpha$in $\mathbb{R}^{N}$ for a given observation, $s$, and the deep critic network predicted the associated state value $V$($s$) or action-value $Q$($s$, $\alpha$), which in turn optimized the PopSAN, in accordance with a chosen DRL method (Fig. 2). Within the PopSAN architecture, the encoder module is responsible for encoding individual dimensions of the observation by mapping them to the activity of distinct neuron populations. During forward propagation, the input populations activate a multi-layer fully-connected SNN. The SNN then produces activity patterns within the output populations. At the end of each set of $T$ timesteps, these patterns of activity are decoded in order to ascertain the associated action dimensions. (as outlined in Algorithm 1).
@@
The current-based leaky-integrate-and-fire (LIF) model of a spiking neuron is employed in constructing the SNN. This model is utilized in constructing the SNN architecture. The dynamics of the LIF neurons are governed by a two-step model, as elaborated in Algorithm 1: i) the integration of presynaptic spikes $o$ into current $c$; and ii) the integration of current $c$ into membrane voltage $v$; $d_{c}$ and $d_{v}$ represent the current and voltage decay factors, respectively. In this particular implementation, a neuron fires a spike when its membrane potential surpasses a predetermined threshold. The hard-reset model was implemented, in which the membrane potential is promptly reset to the resting potential following a spike. The resultant spikes are transmitted to the post-synaptic neurons during the same inference timestep, under the assumption of zero propagation delay. This approach facilitates efficient and synchronized information transmission within the SNN.
@@
Figure 3: RMA consists of two subsystems: the base policy $\pi$ and the adaptation module $\phi$. The RMA training consists of two phases. Training the Base Policy (Phase 1): In the initial phase, the base policy $\pi$ is trained using PopSAN. The system takes the current state $x_{t}$, the previous action $\alpha_{t-1}$, and the environmental factors $e_{t}$ as input. These environmental factors are encoded into a latent extrinsics vector $z_{t}$ using the environmental factor encoder $\mu$. Training the Adaptation Module (Phase 2):In the second phase, the adaptation module $\phi$ is trained to predict the extrinsics $\widehat{z_{t}}$ using past states and actions. This training utilizes supervised learning with on-policy data. The adaptation module learns to capture the relationship between the state-action history and the corresponding extrinsics. By training the base policy and the adaptation module in these two phases, RMA can learn and adapt to the environment more effectively.
Next, we combined the advantages of SNN with the advanced RMA algorithm. Figure 3 shows that the RMA system consists of two interconnected subsystems: the base policy $\pi$ and the adaptation module $\phi$. These subsystems work together to support continuous adaptation in various environmental conditions, ensuring seamless online operations. The base policy is trained using reinforcement learning in simulation, leveraging privileged information about the environment configuration $e_{t}$, such as friction, payload, and other factors. By utilizing the vector $e_{t}$, the base policy can adapt effectively to the unique characteristics of the environment.
@@
Unfortunately, implementing this policy directly is not feasible in reality because $e_{t}$ is unavailable. To address this challenge, we must estimate the extrinsics during runtime, a task performed by the adaptation module $\phi$. The actual movement of robot joints deviates from the intended movement due to extrinsic influences. Instead of depending on privileged information, we can use the agent’s recent state history to estimate the extrinsics vector. The purpose of $\phi$ is to estimate the extrinsics vector $z_{t}$ based solely on the recent state and action history of the robot, without direct access to $e_{t}$.
@@
In addition, we have successfully combined SNN with AMP and achieved similar performance to ANN on legged robots. Figure 4 provides a schematic overview of the system. The motion dataset $M$ consists of a collection of reference motions, where each motion $m^{i}$ $=$ $\widehat{q}_{t}^{i}$ is represented as a sequence of poses $\widehat{q}_{t}^{i}$. The motion clips can be obtained from a variety of sources, including motion capture (mocap) recordings of real-life actors or artist-authored keyframe animations. The simulated robot’s movement is governed by a policy $\pi(\alpha_{t}|s_{t},g)$ that links the character’s state $s_{t}$ and a given goal $g$ to a distribution of actions $\alpha_{t}$. The policy generates actions that determine the desired target positions for proportional-derivative (PD) controllers at each joint of the robot. The controllers then generate control forces to propel the robot’s motion according to specified target positions. The goal $g$ defines a task reward function $r_{t}^{G}=r^{G}(s_{t},\alpha_{t},s_{t+1},g)$ that outlines the high-level objectives the robot needs to achieve. The style objective $r_{t}^{S}=r^{S}(s_{t},s_{t+1})$ is determined by an adversarial discriminator, which provides an a priori estimate of the naturalness or style of a motion, independent of the task at hand. By doing this, the style objective encourages the policy to produce movements that closely mirror the behaviors seen in the dataset.
@@ III-B Training of Legged Robot using SNN on Isaacgym
The SNN parameters are updated using extended spatiotemporal backpropagation as introduced in [39]. We utilized the rectangular function $z(v)$, as defined in [40], to estimate the spike’s gradient. The gradients of the loss with respect to the parameters of the SNN for each layer $k$ are computed by aggregating the gradients backpropagated from all timesteps:
@@ IV Experiments
The objectives of our experiments are the followings: i) To validate the feasibility of SNNs on robots operating in environments with high latitudes and intricate dynamics models. ii) Verify the benefits of SNNs over ANNs in terms of ultra-high flatness control. iii) Identify rewards that SNNs excel on rl tasks and make improvements on poorly performing rewards. We assessed our approach using the Isaac Gym platform, which offers precise physics simulation and a versatile, customizable framework for robot models and environments.
We primarily evaluated the performance of the following robots: A1, Cassie [41], and MIT Humanoid [42] where each SNN is trained within $1,500,000$ iterations until convergence. The ultimate assessment involves utilizing visual representations and metrics, including base linear velocity in the x and y axes, angular velocity in the yaw axis, mean episode length, and position torque are used for evaluation.
@@ IV-A Simulation Setup
In order to establish a varied array of environments, we incorporated multiple URDF files from recent studies. The files contain various models, including A1, Anymal-b, Anymal-c, Cassie, MIT Humanoid, and others. Once imported, we utilize the functionalities of the Isaac Gym simulator to generate novel environments for each of these models. To introduce diversity, we make use of the built-in fractal terrain generator offered by the simulator. This terrain generator facilitates the creation of diverse types of terrain, encompassing flat terrains as well as rugged terrains with a variety of topographical characteristics. The policy functions at a control frequency of up to 500Hz, attributed to our SNN-based approach. This capability enables the attainment of high-frequency control, leading to prompt and precise adjustments to the system. In Table I listed a comprehensive list of environmental variations and their respective ranges is presented.
@@ IV-B Performances of High Frequency Control using SNNs
We conducted tests on the robots mentioned above specifically for linear and angular velocity tracking tasks. In these tests, we considered tracking velocity (linear and angular) as the primary positive rewards. Additionally, we incorporated penalties for base height that is too low, excessive acceleration, and instances where the robot falls, etc. For A1, we conducted training and testing in several terrain environments, including pyramid stairs like terraces (upstairs and downstairs), pyramids with sloping surfaces, hilly terrains, terrains with discrete obstacles, and terrains covered with stepping stones. On the other hand, Cassie is solely trained in a trapezoidal pyramid environment and MIT Humanoid in a plain terrain.
@@
To explore the benefits of SNNs in high-frequency control scenarios, we intentionally increased the simulation environment’s time step (dt) to 2.5 times the default ANNs task, achieving a frequency of 500Hz. Our goal is to determine if SNN outperforms ANN in real-world environments that demand high-frequency control by comparing their performance.
@@
The investigation is motivated by the limitations robots face when using ANN control algorithms due to energy constraints. As a result, these robots can typically only achieve a policy inference frequency of 100Hz, significantly lagging behind the motors’ execution frequency. Instead, SNN has the potential to improve policy inference quality and provide real-time control performance by leveraging its energy-efficient nature and deployment on specialized processors. If SNN can match or outperform ANN in high-frequency control, it would demonstrate SNN’s superiority in real-world environments.
In Figure 5, the effectiveness of our method under high-frequency control is clearly demonstrated. The A1 robot demonstrates impressive performance in tracking velocity x, accurately following the desired trajectory. Considering the impact of complex terrain conditions, all velocities varied within a margin of $17\%$ around the designated value. Although occasional spikes occur, indicating moments when the robot deviates from the existing policy framework, these instances can be attributed to the robot’s inclination to explore alternative approaches when faced with challenging terrain. This emphasizes the robustness of our method in tackling challenging terrains.